OpenStudy (anonymous):

A cone has a radius of 5 inches and a height of 18 inches. What is the exact volume of the cone? A. 540π in3 B. 450π in3 C. 300π in3 D. 150π in3

OpenStudy (diamondboy):

Hi, do u have a formula?

OpenStudy (anonymous):

i dont know it

OpenStudy (diamondboy):



OpenStudy (diamondboy):

r= radius and h= height

OpenStudy (diamondboy):

can u do it now?

OpenStudy (anonymous):

i got 471???

OpenStudy (diamondboy):

wait

OpenStudy (diamondboy):

I got 450 pi

OpenStudy (anonymous):

\[\frac{1}{3} \pi h r^2\text{/.}\, \{r\to 5,h\to 18\}=150 \pi \]/. replaces r and h with 5 and 18 respectively.

OpenStudy (diamondboy):

yep 150 pi sorry